# FLAP: Fully-controllable Audio-driven Portrait Video Generation through 3D head conditioned diffusion model

### v1

@@ 5.3 Ablation Study
We conduct an ablation study to prove the effectiveness of our proposed Progressively Focused Training (PFT) scheme. As is shown in the Figure 4 (the second and third left column), although FLAP can be trained to fit the source pose and expression well with out PFT scheme, its head pose is heavily relied on signals leaked from expression coefficients and fail to decouple them. Once we set its expression coefficients idle, the generated head pose is decreased in range. By contrast, when PFT is applied, the expression change will not affect head pose and achieve disentanglement. In addition, when the model only completes the head motion training, it can generate head motion according to pose source, but is not capable of modifying the expression of reference image. In the last column, when the model completes expression training stage, it’s capable of generating head pose and expression from different source. This decoupling of head pose and facial expression allows for independent control and shows the effectiveness of proposed PFT.
@@ 5.4 Effectiveness of Audio-to-FLAME module
Our Audio-to-FLAME module is capable of binding the facial expression to audio, making it reference image irrelevant. As shown in Figure 5, Hallo struggles to generate appropriate expressions since it is confused by the smiling face in reference image and angry tone in the audio. Therefore, it generates a mixed and unnatural face. On the contrary, our model generate expressions solely relies on audio though our Audio-to-FLAME module, and thus generate accurate and natural expressions.
@@ 6 Conclusion
In this paper, we present FLAP, a novel approach to zero-shot talking-head avatar generation that integrates explicit 3D head pose and facial expression coefficients into a diffusion-based framework. By enabling independent control over head movements and facial expressions, FLAP significantly improves both the naturalness and controllability of generated portrait videos. Our model demonstrates a nuanced understanding of 3D space, offering flexibility in controlling head movements and generating lifelike expressions from diverse sources, including audio and user-specified coefficients. Extensive experiments show that FLAP outperforms existing methods in terms of visual quality, flexibility and controllability, providing a promising solution for real-world applications.
@@
Appendix A Appendix
A.1 More Results for Flexibility Analysis
Figure 1 demonstrates the visual results of integrating style controlled Audio-DVP with FLAP. As shown in Figure 1 With audio-dvp, the style of mouth while speaking is changed, a greater degree of jaw opening is generated. Thus, by combining with various existing methods, we can generate a range of speaking styles that we desire.
Figure 1: Visual result of integrating talking style from (Zhang et al., 2023b) into Audio-DVP(Wen et al., 2020) with FLAP. The first row demonstrate original FLAP, while the second row shows the result of utilizing Audio-DVP + style control as a replacement of Audio-to-FLAME module.
Figure 2: Visual results of FLAP framework cooperating with other methods. The first and second rows are head pose and expression source images respectively. The other rows are id reference and generated results.
@@ A.2 Flexibility Analysis
FLAP exhibits a high level of flexibility as it can be integrated with existing models by using their feature vectors as diffusion condition. As an example, As an example, EDtalk (Tan et al., 2025) proposed a method to extract implicit representations of facial expressions, head pose, lip pose and identity information from a single portrait image. Therefore, we can replace our 3D head head conditions in FLAP with feature vectors provided by EDtalk. Figure 2 demonstrates the visual results of retraining FLAP with expression and head pose conditions from EDtalk. The generated results have the same pose as the pose reference image, and the same expression as the expression reference image. Pose and expression features extracted by EDtalk(Tan et al., 2025) are used as intermediate representations.
Further, the replaceable audio-to-FLAME module offers flexibility. As an example, retraining Audio-DVP (Wen et al., 2020) and add style embedding following (Zhang et al., 2023b) can change the style of talking. Visual results are in the supplementary materials.
Figure 3: More portrait animation results. Reference images are from Loopy(Jiang et al., 2024), EchoMimic (Chen et al., 2024b), Follow-Your-Emoji (Ma et al., 2024b) and EMO (Tian et al., 2024) and HDTF dataset.
A.3 More User Controlled Animation Results
We provide more animation results in Figure 3. In the last row, we animate a portrait with audio signal only and set her head pose fixed. In the rest of the rows, we controlled the characters to turn their heads, nod, look upwards, and tilt their heads. In the meantime, they maintain speaking according to the same audio input of the last row. We highlight that all these head pose controls are done by inputting the head rotation angles, no additional video signal is involved.
Figure 4: Our methods avoids the mean face and mean head problem of vasa (Xu et al., 2024b)
A.4 Comparison with VASA
Since vasa (Xu et al., 2024b) is not a open source project, we can only give comparison based on images in its paper. Here we find vasa official demo image suffers severe mean face problem, losing its hair and face feature of the reference image. Vasa model seems tend to generate a face of a average man and fail to capture the detailed feature of reference image. From our observation, this frequently happened in most FaceVid2Vid(Wang et al., 2021) backboned models. In fact, vasa uses a retrained FaceVid2Vid as its image encoder and decoder, which limit its expressiveness. Figure 4 shows our method can successfully avoid this mean face problem.

### v2

@@ 5.5.2 Ablation Study
We conduct an ablation study to demonstrate the effectiveness of our proposed Progressively Focused Training (PFT) scheme. As is shown in the Figure 7 (the 2nd and 3rd left column), although FLAP can be trained to fit the source pose and expression well on video driven mode without PFT scheme, its head pose is heavily relied on signals leaked from expression coefficients, indicating failure in decoupling. To confirm this assumption, in the 4rd column, we set its expression coefficients idle, and get a result with decrease of head movement range. By contrast, in 5th column, when PFT is applied, the expression change will not affect head pose and achieve full disentanglement. In addition, when the model only completes the head motion training, it can only generate head motion aligning with pose source, but is not capable of modifying the expression of reference image (6th column). In the last column, when the model completes expression training stage, it’s capable of generating both head pose and expression from two different source. This decoupling of head pose and facial expression allows for independent control and shows the effectiveness of proposed PFT. Moreover, the proposed PFT scheme can be generalized to other conditions during training. We test this scheme useing feature vectors from EDTalker (Tan et al., 2025) in supplementary.
Figure 8: Visual result of our method and Hallo2 (Cui et al., 2024) animating a smiling portrait with angry tone.
@@ 5.6 Effectiveness of Audio-to-FLAME module
Our Audio-to-FLAME module is capable of binding the facial expression to audio, making all expressions irrelevant to reference image. As shown in Figure 8, Hallo2 struggles to generate appropriate expressions since it is confused by the smiling face in reference image and angry tone in the audio. Therefore, it generates a mixed and unnatural face. On the contrary, our model generate expressions solely relies on audio though our Audio-to-FLAME module, and thus generate accurate and natural expressions.
@@ 6 Conclusion
In this paper, we present FLAP, a novel approach for zero-shot talking-head avatar generation that integrates explicit 3DMM coefficients into a diffusion-based framework. By enabling independent control over head movements and facial expressions with PFT, FLAP significantly improves both naturalness and controllability of generated portrait videos. Our model demonstrates a nuanced understanding of 3D space, offering flexibility in controlling head movements and generating lifelike expressions from diverse sources. Extensive experiments show that FLAP outperforms existing methods in terms of visual quality lip synchronization and controllability, providing a promising solution for real-world applications.
@@
A Appendix
Figure 1: Additional Results on Multi Pose Generation
A.1 Additional Results on Multi Pose Generation
Figure 1 provides more results of multi-pose and multi-expression generation of the same identity. The range of head rotation angle is pre-defined by a 3-dim vector representing [x,y,z] angle in 3 DoF as control signal from user.
@@ A.2 Flexibility Analysis
A.2.1 Portability of FLAME Coefficients
As discussed in Sec4.4, FLAME coefficients is portable, making our audio-to-Flame module replaceable by any FLAME coefficients generating methods such as EmoTalk (Peng et al., 2023), EMOTE (Daněček et al., 2023), DiffPoseTalk (Sun et al., 2024). For demonstration, we implemented a talking style controllable FLAME talker by integrating talking style control mechanism from (Zhang et al., 2023b) into Audio-DVP (Wen et al., 2020). Figure 2 demonstrates the visual results of integrating this style controlled Audio-DVP into FLAP. As shown in Figure 2, with this audio-dvp model, the style of mouth while speaking is changed, compared with original Audio-2-FLAME based FLAP, a greater degree of jaw opening is generated making it easier for lip reading. Thus, by combining with various existing methods, we can generate a range of speaking styles that we desire.
Figure 2: Visual result of integrating talking style from (Zhang et al., 2023b) into Audio-DVP(Wen et al., 2020) and cooperate with FLAP. The first row demonstrate original FLAP, while the second row shows the result of utilizing Audio-DVP + style control as a replacement of Audio-to-FLAME module.
Figure 3: Visual results of FLAP framework cooperating with PD-FGC (Wang et al., 2023) by retraining FLAP using feature vectors from PD-FGC as conditions. We perform pose alignment of driving image using pose vector from PD-FGC.
Figure 4: Visual results of FLAP framework cooperating with EDtalk (Tan et al., 2025). The first and second columns are head pose and expression references respectively. The first row shows the id reference. The others are generated results with mixed pose, expression and id of corresponding reference.
A.2.2 Using Feature Vectors as Conditions
@@
FLAP exhibits a high level of flexibility as it can be integrated with existing models by using their feature vectors as diffusion condition. For example, EDtalk (Tan et al., 2025) and PD-FGC (Wang et al., 2023) proposed methods to extract implicit representations of facial expressions, head pose, lip pose and identity information from a single portrait image. Therefore, we can replace our 3D head head conditions in FLAP with feature vectors provided by EDtalk or PD-FGC. Figure 3 demonstrates the results of pose alignment using pose vector from PD-FGC. Figure 4 shows the results of independent control of pose and expression using pose and expression feature vector from EDTalk. Although we achieve controllable using these feature vectors, its only yield results with less naturalness and expressiveness than our original model.
A.3 Additional Effectiveness Analysis of PFT Scheme
To test the generalize ability of our Progressively Focused Training (PFT) scheme, we retrained our FLAP frame work using feature vectors from EDtalk (Tan et al., 2025) as conditions for denoising. From our observation, the feature vector of EDtalk is not fully disentangled since its generator is trained to extract desired expression and pose information from these feature. Thus, PFT is necessary in training process to make sure disentanglement of pose and expressions. As shown in Figure 4, this retrained FLAP generates results with their pose aligned with pose reference image and expression aligned with expression reference image shown in the first two columns. This not only shows the generalize ability of PFT, but also demonstrate the flexibility of our model structure.
Figure 5: Hand artifact of existing models.
Figure 6: Detection of a small part of the fingertip.
A.4 Hands Filtering in Data Pre-Processing
Filtering frames with hands in training data is crucial for the final generation results. As shown in Figure 6, existing open sourced models like Echomimic and Aniportrait is trained on unclean dataset in which hands often appears. To train a model that will not generate artifact of hands or fingers, we utilize Grounding Dino to detect hands and fingers every 3 frames in a video. As shown in Figure 6, even a little spot of finger can be detected and cut out from the video. Thus, we successully trained a clean model on clean dataset.
A.5 More Animation Results
We provide more animation results in the next page, to demonstrate the generalization ability our our model and also to illustrate their visual quality. Please refer to Figure 7 in the NEXT PAGE to see diverse expressions, poses and styles of our generated results.
Figure 7: More Animation Results